# Boaz Moerman

### v2

@@ 1. Introduction
Over the last few decades there has been a focused effort to count the number of rational points on algebraic varieties and special subsets thereof. Manin’s conjecture is a major source of inspiration for many of these developments, which gives an asymptotic formula for the number of rational points of bounded height on rationally connected varieties. This conjecture has been refined over several decades [FMT89, BaMa90, Pey95, BaTs98, LeSeKuSh22], and it has been established for several families of varieties (see e.g. [BaTs96, GoMaOh08, Woo24]).
In [PSTVA21], Pieropan, Smeets, Tanimoto and Várilly-Alvarado proposed an asymptotic formula for the number of Campana points of bounded height, extending Manin’s conjecture. Campana points are rational points which intersect a given collection of boundary divisors $D_{1},\dots,D_{n}$ with high multiplicity, and they arise naturally as powerful solutions of Diophantine equations. Their conjecture has been established in various special cases, as we will discuss below.
Besides Campana points, various other related subsets have been studied recently, including weak Campana points, geometric Campana points and (geometric) Darmon points. Several asymptotic counting results have been established for these points [Str21, ShSt24, Ito25, Ara25], in the spirit of Manin’s conjecture. Despite these results, no extension of Manin’s conjecture to these types of points has been proposed thus far.
These various special sets of rational points mentioned are all instances of sets of $\mathcal{M}$-points, which were first introduced in [Moe24]. Many other arithmetically interesting sets of rational points can be viewed as $\mathcal{M}$-points, such as the set of squarefree solutions to a given Diophantine equation. Similarly, the set
@@
is also a set of $\mathcal{M}$-points, for any given homogeneous integer polynomial $f$. Such sets are related to the study of squarefree values of polynomials, which is an enduring and active research area, see for example [Hoo67, Fil92, Poo03, SaWa23].
In this article we introduce Conjecture 1.4, which gives an asymptotic formula for the number of $\mathcal{M}$-points of bounded height, further extending Manin’s conjecture. We show that the conjecture agrees with the prediction given for Campana points in [PSTVA21], with the corrected leading constant formulated in [CLTT24, Conjecture 8.3].
Our conjecture also gives an asymptotic for Campana points of bounded height for more Campana pairs than the conjecture in [PSTVA21]. One major limitation in the applicability of the latter conjecture is that it requires the boundary to be a strict normal crossings divisor. In contrast, our conjecture only assumes that the boundary is a normal crossings divisor after passing to the algebraic closure of the base field. In particular, our conjecture explains the asymptotic formula [Str21] found by Streeter for the number of powerful values of norm forms. His results did not match the asymptotic formula given by (a naive extension) of the conjecture [PSTVA21], but his results agree with our new conjecture, as we will show.
More generally, the weakened hypothesis makes Conjecture 1.4 applicable in much more situations than the previous conjecture. For instance, our conjecture gives a prediction for the number of Campana points of bounded height on a toric variety (with torus-invariant boundary), while the conjecture in [PSTVA21] is only applicable for special toric varieties such as split toric varieties.
Furthermore, our conjecture also predicts an asymptotic for the number of weak Campana points of bounded height. In [PSTVA21], a preliminary investigation was done into such asymptotics, but the authors concluded that a conjecture for them would require a new set of ideas.
Finally, we verify that the conjecture agrees with the various results [Str21, ShSt24, Ito25, Ara25] known on $\mathcal{M}$-points. We will treat Streeter’s result [Str21, Theorem 1.1] on powerful values of norm forms in great detail, to show it agrees with our prediction. In particular, this gives a geometric interpretation for the exponent on the logarithm appearing in the asymptotic, which was not known previously [Str21, Remark 1.5]. In an upcoming work [Moe25toric], we prove this conjecture for split toric varieties over the rational numbers, giving the first general results on counting $\mathcal{M}$-points.
@@ 1.1. Manin’s conjecture and Campana points
for the height on $X$ as defined in [Pey95, §1.3]. We denote the corresponding counting function by
@@
where $A\subset X(K)$ and $B$ is an integer. We recall the most recent version of Manin’s conjecture conjecture, which is given for example in [LeSeKuSh22, Conjecture 1.2].
@@
Manin’s conjecture was first formulated and studied for Fano varieties with an anticanonical height in 1989 and 1990 by Manin, Batyrev, Tschinkel and Franke [FMT89, BaMa90]. Peyre [Pey95] further contributed to the conjecture by giving a conjectural value for the constant $c$. Later, Batyrev and Tschinkel extended the conjecture to heights induced by other divisors. More recently Lehmann, Sengupta and Tanimoto [LeSeKuSh22] have formulated a prediction for the thin set $Z$ that has to be excluded.
@@
is a boundary divisor given by weights $m_{1},\dots,m_{n}\in\mathbb{N}$ and prime divisors $D_{1},\dots,D_{n}$ on $X$. Let $S\subset\Omega_{K}$ be a finite set of places of $K$ including the infinite places and let $\mathcal{X}$ be an $\mathcal{O}_{S}$-integral model of $X$. Let $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$ be the corresponding $\mathcal{O}_{S}$-integral model of $(X,D_{\mathbf{m}})$ and write $(\mathcal{X},\mathcal{D}_{\mathbf{m}})(\mathcal{O}_{S})$ for its set of Campana points over $\mathcal{O}_{S}$. In [PSTVA21], Pieropan, Smeets, Tanimoto and Várilly-Alvarado formulated the following conjecture on the number of Campana points of bounded height.
@@
where $c>0$ is a constant,
@@
Furthermore, if this divisor class is rigid, then $c=c(K,S,(\mathcal{X},\mathcal{D}_{\mathbf{m}}),\mathcal{L})$ is an explicit Tamagawa constant, given in [CLTT24, Conjecture 8.3].
The original conjecture, as given in [PSTVA21], had a different prediction for the leading constant. However, Shute [Shu22] and Streeter [Str21] independently gave counterexamples to this prediction. In [CLTT24], Chow, Loughran, Takloo-Bighash and Tanimoto modified the conjecture by changing the Tamagawa constant to the one cited above, so that Shute’s and Streeter’s results are in agreement with their conjecture. Their constant is defined very similar to the original constant, but it is defined using a different Brauer group.
Conjecture 1.2 has been proven for various Campana pairs. These include results on diagonal hypersurfaces [Val12, BrYa21, Shu21, Shu22, BBKOPW23], vector group compactifications [PSTVA21], norm forms [Str21], biequivariant compactifications of the Heisenberg group [Xia21] and wonderful compactifications [CLTT24]. Furthermore, for the log-anticanonical height, this conjecture has been proven for complete toric varieties [PiSc23, ShSt24] and certain complete intersections therein [PiSc24].
Furthermore, in [Fai23, Fai25] Faisant proves a motivic analogue of the conjecture for Campana points for vector group compactifications and toric varieties.
The assumption in Conjecture 1.2 that $\operatorname{Supp}D_{\mathbf{m}}$ is a strict normal crossings divisor is rather restrictive, as many interesting Campana pairs do not satisfy this hypothesis. For instance, this assumption need not be satisfied if $X$ is a smooth projective toric variety and $\operatorname{Supp}D_{\mathbf{m}}$ is the complement of the open torus. Nevertheless, this assumption is important as Streeter shows in [Str21] that the logarithmic factor appearing in the asymptotic is larger than the conjecture would suggest. In order to avoid this issue, Shute and Streeter introduced the notion of geometric Campana points [Str21, ShSt24], which are a variant on Campana points defined using the geometric components of $D_{\mathbf{m}}$ (rather than those defined over $K$). For these points, they showed that Conjecture 1.2 holds for the geometric Campana points on toric Campana pairs with respect to the log-anticanonical height.
In a similar fashion, analogues of Conjecture 1.2 have been proven for (geometric) Darmon points [ShSt24, Ito25, Ara25].
@@ 1.2. $\mathcal{M}$-points
One thing in common between Campana points, Darmon points and their variants, is that they are defined as rational points satisfying certain conditions on their intersection multiplicities with the boundary components. In this article, we will extend Conjecture 1.2 to the setting of $\mathcal{M}$-points. $\mathcal{M}$-points are a broad generalization of Campana points, Darmon points, as well as their variants mentioned above. They were first introduced by the author in [Moe24]. Like Campana points, $\mathcal{M}$-points are rational points satisfying conditions on their intersection multiplicities with the boundary components.
@@
More generally, we also allow the set $M$ to impose conditions with respect to the different intersection strata. The notion of pair used in this article is more general than the notion considered in [Moe24, Thesis], as we explain in Section 3, where we define pairs and $\mathcal{M}$-points.
@@ 1.3. $\mathcal{M}$-points of bounded height
Now we will formulate a version of Manin’s conjecture for $\mathcal{M}$-points. Let $(X,M)$ be a smooth, proper and rationally connected pair over a number field $K$, as defined in Section 5. For such a pair, we introduce its Picard group $\operatorname{Pic}(X,M)$ along with a natural group homomorphism $\operatorname{pr}^{*}_{M}\colon\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,M)$, its pseudo-effective cone $\overline{\operatorname{Eff}}^{1}(X,M)$ as well as its canonical divisor class $K_{(X,M)}\in\operatorname{Pic}(X,M)$. Using these notions, we formulate our conjecture exactly analogous to Manin’s conjecture. Fix a finite set $S$ of places of $K$, including all infinite places and the places that ramify in the splitting field of $(X,M)$. For an integral model $(\mathcal{X},\mathcal{M})$ of $(X,M)$ over $\mathcal{O}_{S}$, we denote the set of $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$ over $\mathcal{O}_{S}$ by $(\mathcal{X},\mathcal{M})(\mathcal{O}_{S})\subset X(K)$. As before, we let $L$ be a big and nef divisor class with adelic metrization $\mathcal{L}$.
@@
Let $(X,M)$ be a smooth proper pair over a number field $K$ such that $(X,M)$ is rationally connected, and let $(\mathcal{X},\mathcal{M})$ be an integral model of $(X,M)$ over $\mathcal{O}_{S}$. Assume furthermore that $(\mathcal{X},\mathcal{M})(\mathcal{O}_{S})\subset X(K)$ is Zariski dense in $X$. Then there exists a thin set $Z\subset X(K)$ such that
@@
This conjecture is an extension of the conjecture [Thesis, Conjecture 1.2.2] presented in the author’s PhD thesis, where only split pairs were considered and no prediction for the leading constant was given.
@@
The assumption that $(X,M)$ is smooth is a much more relaxed assumption that the assumption in Conjecture 1.2 that $D_{\mathbf{m}}$ has strict normal crossings support. For instance, the conjecture can be applied to any smooth toric variety, where the boundary is taken to be the complement of the torus.
@@
The assumption on the fundamental group is satisfied whenever the complement $U$ of the boundary divisors has an abelian fundamental group (which is satisfied in all cases for which the conjecture is known). In personal communication, Tim Santens has suggested a Campana pair for which we expect that the prediction in Conjecture 1.4 for the constant fails due to the existance of a nonabelian cover. This assumption was also suggested to be necessary in [CLTT24] for the same reason, see [CLTT24, Remark 8.4].
@@
The condition on the effective cone of $X$ is quite weak, as it is satisfied if $X$ is a Mori dream space, such as when $X$ is a Fano variety [BCHM10, Corollary 1.3.2] or a toric variety.
Theorem 1.7 has a variety of conferences. Firstly, it shows that Conjecture 1.2 is a special case of Conjecture 1.4, provided that the relevant pair $(X,M)$ is rationally connected. This rational connectedness follows from a conjecture of Campana [Cam11, Conjecture 9.10], which we will extend in Section 6 to other pairs. In Section 9, we will explain why the Tamagawa constant in our conjecture agrees with the Tamagawa constant in [CLTT24].
@@
As described in [Moe24, §8], Darmon points correspond to integral points on the corresponding root stack. Conjecture 1.4 therefore gives a prediction for the asymptotic number of integral points of bounded height on a root stack $(\mathcal{X},\sqrt[\mathbf{m}]{\mathcal{D}})$, where the height is induced by any metrized big and nef line bundle on $X$. This is related to the conjecture by Ellenberg, Satriano and Zureick-Brown on the number of rational points of bounded height on stacks [ESZB23, Conjecture 4.14] and its generalization by Darda and Yasuda [DaYa24-general, Conjecture 9.16]. Their conjectures use different heights however: the heights we consider are what Darda and Yasuda call unstable heights, while their conjecture uses stable heights instead.
@@
In contrast to Conjecture 1.1 and Conjecture 1.2, we do not assume that the set of $\mathcal{M}$-points in Conjecture 1.4 is not thin, but we only require that it is Zariski dense. The reason for this is that [Moe25toric] implies the conclusion of Conjecture 1.4 for every smooth proper toric pair, even though there are many such pairs $(X,M)$ for which the set of $\mathcal{M}$-points is thin by [Moe24, Theorem 6.5]. Furthermore, there are no known examples of rationally connected varieties with a thin, but nonempty, set of rational points. Any such example would contradict an open conjecture by Colliot-Thélène [Ser16, Conjecture 3.5.8], [CoT88].
@@
In Conjecture 1.2, the assumption is made that the integral model $\mathcal{X}$ is regular, which we do not assume in Conjecture 1.4. We suspect that this assumption is superfluous, supported by the results by Santens [San23-2, Theorem 1.1] providing an analogue of Manin’s conjecture for integral points on toric varieties with respect to any flat integral model.

### v3

# Manin’s conjecture for $\mathcal{M}$-points

@@ 1. Introduction
Over the last few decades there has been a focused effort to count the number of rational points on algebraic varieties and special subsets thereof. Manin’s conjecture is a major source of inspiration for many of these developments, which gives an asymptotic formula for the number of rational points of bounded height on rationally connected varieties. This conjecture has been refined over several decades [21, 3, 34, 6, 27], and it has been established for several families of varieties (see e.g. [5, 22, 52]).
In [37], Pieropan, Smeets, Tanimoto and Várilly-Alvarado proposed an asymptotic formula for the number of Campana points of bounded height, extending Manin’s conjecture. Campana points are rational points which intersect a given collection of boundary divisors $D_{1},\dots,D_{n}$ with high multiplicity, and they arise naturally as powerful solutions of Diophantine equations. Their conjecture has been established in various special cases, as we will discuss below.
Besides Campana points, various other related subsets have been studied recently, including weak Campana points, geometric Campana points and (geometric) Darmon points. Several asymptotic counting results have been established for these points [49, 45, 25, 1], in the spirit of Manin’s conjecture. Despite these results, no extension of Manin’s conjecture to these types of points has been proposed thus far.
These various special sets of rational points mentioned are all instances of sets of $\mathcal{M}$-points, which were first introduced in [30]. Many other arithmetically interesting sets of rational points can be viewed as $\mathcal{M}$-points, such as the set of squarefree solutions to a given Diophantine equation. Similarly, the set
@@
is also a set of $\mathcal{M}$-points, for any given homogeneous integer polynomial $f$. Such sets are related to the study of squarefree values of polynomials, which is an enduring and active research area, see for example [24, 20, 38, 42].
In this article we introduce Conjecture 1.4, which gives an asymptotic formula for the number of $\mathcal{M}$-points of bounded height, further extending Manin’s conjecture. We show that the conjecture agrees with the prediction given for Campana points in [37], with the corrected leading constant formulated in [13, Conjecture 8.3].
Our conjecture also gives an asymptotic for Campana points of bounded height for more Campana pairs than the conjecture in [37]. One major limitation in the applicability of the latter conjecture is that it requires the boundary to be a strict normal crossings divisor. In contrast, our conjecture only assumes that the boundary is a normal crossings divisor after passing to the algebraic closure of the base field. In particular, our conjecture explains the asymptotic formula [49] found by Streeter for the number of powerful values of norm forms. His results did not match the asymptotic formula given by (a naive extension) of the conjecture [37], but his results agree with our new conjecture, as we will show.
More generally, the weakened hypothesis makes Conjecture 1.4 applicable in much more situations than the previous conjecture. For instance, our conjecture gives a prediction for the number of Campana points of bounded height on a toric variety (with torus-invariant boundary), while the conjecture in [37] is only applicable for special toric varieties such as split toric varieties.
Furthermore, our conjecture also predicts an asymptotic for the number of weak Campana points of bounded height. In [37], a preliminary investigation was done into such asymptotics, but the authors concluded that a conjecture for them would require a new set of ideas.
Finally, we verify that the conjecture agrees with the various results [49, 45, 25, 1] known on $\mathcal{M}$-points. We will treat Streeter’s result [49, Theorem 1.1] on powerful values of norm forms in great detail, to show it agrees with our prediction. In particular, this gives a geometric interpretation for the exponent on the logarithm appearing in the asymptotic, which was not known previously [49, Remark 1.5]. In another work [32], we prove this conjecture for split toric varieties over the rational numbers, giving the first general results on counting $\mathcal{M}$-points.
@@ 1.1. Manin’s conjecture and Campana points
for the height on $X$ as defined in [34, §1.3]. We denote the corresponding counting function by
@@
where $A\subset X(K)$ and $B$ is an integer. We recall the most recent version of Manin’s conjecture conjecture, which is given for example in [27, Conjecture 1.2].
@@
Manin’s conjecture was first formulated and studied for Fano varieties with an anticanonical height in 1989 and 1990 by Manin, Batyrev, Tschinkel and Franke [21, 3]. Peyre [34] further contributed to the conjecture by giving a conjectural value for the constant $c$. Later, Batyrev and Tschinkel extended the conjecture to heights induced by other divisors. More recently Lehmann, Sengupta and Tanimoto [27, Conjecture 5.2] have formulated a prediction for the thin set $Z$ that has to be excluded.
@@
is a boundary divisor given by weights $m_{1},\dots,m_{n}\in\mathbb{N}$ and prime divisors $D_{1},\dots,D_{n}$ on $X$. Let $S\subset\Omega_{K}$ be a finite set of places of $K$ including the infinite places and let $\mathcal{X}$ be an $\mathcal{O}_{S}$-integral model of $X$. Let $(\mathcal{X},\mathcal{D}_{\mathbf{m}})$ be the corresponding $\mathcal{O}_{S}$-integral model of $(X,D_{\mathbf{m}})$ and write $(\mathcal{X},\mathcal{D}_{\mathbf{m}})(\mathcal{O}_{S})$ for its set of Campana points over $\mathcal{O}_{S}$. In [37], Pieropan, Smeets, Tanimoto and Várilly-Alvarado formulated the following conjecture on the number of Campana points of bounded height.
@@
where $c=c(K,S,(\mathcal{X},\mathcal{D}_{\mathbf{m}}),\mathcal{L})>0$ is a constant,
@@
Furthermore, if this divisor class is rigid, then $c$ is an explicit Tamagawa constant, given in [13, Conjecture 8.3].
The original conjecture, as given in [37], had a different prediction for the leading constant. However, Shute [47] and Streeter [49] independently gave counterexamples to this prediction. In [13], Chow, Loughran, Takloo-Bighash and Tanimoto modified the conjecture by changing the Tamagawa constant to the one cited above, so that Shute’s and Streeter’s results are in agreement with their conjecture. Their constant is defined very similarly to the original constant, but it is defined using a different Brauer group which we recall in Section 7.
Conjecture 1.2 has been proven for various Campana pairs. These include results on diagonal hypersurfaces [50, 8, 46, 47, 2], vector group compactifications [37], norm forms [49], biequivariant compactifications of the Heisenberg group [53] and wonderful compactifications [13]. Furthermore, for the log-anticanonical height, this conjecture has been proven for complete toric varieties [35, 45] and certain complete intersections therein [36].
Furthermore, in [18, 19] Faisant proves a motivic analogue of the conjecture for Campana points for vector group compactifications and toric varieties.
The assumption in Conjecture 1.2 that $\operatorname{Supp}D_{\mathbf{m}}$ is a strict normal crossings divisor is rather restrictive, as many interesting Campana pairs do not satisfy this hypothesis. For instance, this assumption need not be satisfied if $X$ is a smooth projective toric variety and $\operatorname{Supp}D_{\mathbf{m}}$ is the complement of the open torus. Nevertheless, this assumption is important as Streeter shows in [49] that the logarithmic factor appearing in the asymptotic is larger than the conjecture would suggest. In order to avoid this issue, Shute and Streeter introduced the notion of geometric Campana points [49, 45], which are a variant on Campana points defined using the geometric components of $D_{\mathbf{m}}$ (rather than those defined over $K$). They showed that Conjecture 1.2 holds for the geometric Campana points on toric Campana pairs with respect to the log-anticanonical height.
In a similar fashion, analogues of Conjecture 1.2 have been proven for (geometric) Darmon points [45, 25, 1].
@@ 1.2. $\mathcal{M}$-points
One thing in common between Campana points, Darmon points and their variants, is that they are defined as rational points satisfying certain conditions on their intersection multiplicities with the boundary components. In this article, we will extend Conjecture 1.2 to the setting of $\mathcal{M}$-points. $\mathcal{M}$-points are a broad generalization of Campana points, Darmon points, as well as their variants mentioned above. They were first introduced by the author in [30]. Like Campana points, $\mathcal{M}$-points are rational points satisfying conditions on their intersection multiplicities with the boundary components.
@@
More generally, we also allow the set $M$ to impose conditions with respect to the different intersection strata, as we discuss in Section 3. The notion of pair used in this article is more general than the notion considered in [30, 31], as we explain in Section 3, where we define pairs and $\mathcal{M}$-points.
@@ 1.3. $\mathcal{M}$-points of bounded height
Now we will formulate a version of Manin’s conjecture for $\mathcal{M}$-points. Let $(X,M)$ be a smooth, proper and rationally connected pair over a number field $K$, as defined in Section 5. For such a pair, we introduce its Picard group $\operatorname{Pic}(X,M)$ along with a natural group homomorphism $\operatorname{pr}^{*}_{M}\colon\operatorname{Pic}(X)\rightarrow\operatorname{Pic}(X,M)$, its pseudo-effective cone $\overline{\operatorname{Eff}}^{1}(X,M)$ as well as its canonical divisor class $K_{(X,M)}\in\operatorname{Pic}(X,M)$. Using these notions, we formulate our conjecture in exact analogy to Manin’s conjecture. Fix a finite set $S$ of places of $K$, including all infinite places and the places that ramify in the splitting field of $(X,M)$. For an integral model $(\mathcal{X},\mathcal{M})$ of $(X,M)$ over $\mathcal{O}_{S}$, we denote the set of $\mathcal{M}$-points on $(\mathcal{X},\mathcal{M})$ over $\mathcal{O}_{S}$ by $(\mathcal{X},\mathcal{M})(\mathcal{O}_{S})\subset X(K)$. As before, we let $L$ be a big and nef divisor class with adelic metrization $\mathcal{L}$.
@@
Let $(X,M)$ be a smooth proper pair over a number field $K$ such that $(X,M)$ is rationally connected, and let $(\mathcal{X},\mathcal{M})$ be an integral model of $(X,M)$ over $\mathcal{O}_{S}$. Assume furthermore that $(\mathcal{X},\mathcal{M})(\mathcal{O}_{S})\subset X(K)$ is Zariski-dense in $X$. Then there exists a thin set $Z\subset X(K)$ such that
@@
This conjecture is an extension of the conjecture [31, Conjecture 1.2.2] presented in the author’s PhD thesis, where only split pairs were considered and no prediction for the leading constant was given.
@@
The assumption that $(X,M)$ is smooth is a much more relaxed assumption than the assumption in Conjecture 1.2 that $D_{\mathbf{m}}$ has strict normal crossings support. For instance, the conjecture can be applied to any smooth toric variety, where the boundary is taken to be the complement of the torus.
@@
The assumption on the fundamental group is satisfied whenever the complement $U$ of the boundary divisors has an abelian fundamental group (which is satisfied in all cases for which the conjecture is known). In personal communication, Tim Santens has suggested a Campana pair for which we expect that the prediction in Conjecture 1.4 for the constant fails due to the existence of a nonabelian cover. This assumption was also suggested to be necessary in [13] for the same reason, see [13, Remark 8.4].
@@
The condition on the effective cone of $X$ is quite weak, as it is satisfied if $X$ is a Mori dream space, such as when $X$ is a Fano variety [7, Corollary 1.3.2] or a toric variety.
Theorem 1.7 has a variety of consequences. Firstly, it shows that Conjecture 1.2 is a special case of Conjecture 1.4, provided that the relevant pair $(X,M)$ is rationally connected. This rational connectedness follows from a conjecture of Campana [10, Conjecture 9.10], which we will extend in Section 6 to other pairs. In Section 9, we will explain why the Tamagawa constant in our conjecture agrees with the Tamagawa constant in [13].
@@
As described in [30, §8], Darmon points correspond to integral points on the corresponding root stack. Conjecture 1.4 therefore gives a prediction for the asymptotic number of integral points of bounded height on a root stack $(\mathcal{X},\sqrt[\mathbf{m}]{\mathcal{D}})$, where the height is induced by any metrized big and nef line bundle on $X$. This is related to the conjecture by Ellenberg, Satriano and Zureick-Brown on the number of rational points of bounded height on stacks [17, Conjecture 4.14] and its generalization by Darda and Yasuda [16, Conjecture 9.16]. Their conjectures use different heights however: the heights we consider are what Darda and Yasuda call unstable heights, while their conjecture uses stable heights instead.
@@
In contrast to Conjecture 1.1 and Conjecture 1.2, we do not assume that the set of $\mathcal{M}$-points in Conjecture 1.4 is not thin, but we only require that it is Zariski-dense. The reason for this is that by [32, Theorem 1.1], Conjecture 1.4 holds for every smooth proper toric pair over $\mathbb{Q}$, even though there are many such pairs $(X,M)$ for which the set of $\mathcal{M}$-points is thin by [30, Theorem 6.5]. Furthermore, there are no known examples of rationally connected varieties with a thin, but nonempty, set of rational points. Any such example would contradict an open conjecture by Colliot-Thélène [44, Conjecture 3.5.8], [15].
@@
In Conjecture 1.2, the assumption is made that the integral model $\mathcal{X}$ is regular, which we do not assume in Conjecture 1.4. We suspect that this assumption is superfluous, supported by the results by Santens [43, Theorem 1.1] providing an analogue of Manin’s conjecture for integral points on toric varieties with respect to any flat integral model.